# PerTouch: VLM-Driven Agent for Personalized and Semantic Image Retouching

### v1

@@ Abstract
Image retouching aims to enhance visual quality while aligning with users’ personalized aesthetic preferences. To address the challenge of balancing controllability and subjectivity, we propose a unified diffusion-based image retouching framework called PerTouch. Our method supports semantic-level image retouching while maintaining global aesthetics. Using parameter maps containing attribute values in specific semantic regions as input, PerTouch constructs an explicit parameter-to-image mapping for fine-grained image retouching. To improve semantic boundary perception, we introduce semantic replacement and parameter perturbation mechanisms in the training process. To connect natural language instructions with visual control, we develop a VLM-driven agent that can handle both strong and weak user instructions. Equipped with mechanisms of feedback-driven rethinking and scene-aware memory, PerTouch better aligns with user intent and captures long-term preferences. Extensive experiments demonstrate each component’s effectiveness and the superior performance of PerTouch in personalized image retouching. Code is available at: https://github.com/Auroral703/PerTouch.
@@ 3.2 Architecture
We observe that directly injecting the parameter map into the network leads to over-reliance on the information it encodes, which is contrary to our objective. Rather than enforcing strict adherence to the parameter map, we aim to use it as a soft hint, allowing the diffusion prior to play a central role in generating aesthetically pleasing results. To mitigate the model’s over-reliance on the parameter map, we introduce two mechanisms.
First, we find that the model struggles to perceive region boundaries based on the injected parameter map, often defaulting to global retouching. This is because the input map contains only coarse semantic scores, which are not spatially continuous across regions, while real images often exhibit spatially coherent color transitions. This discrepancy makes it difficult for the model to correlate the parameter map with the image structure. To address this, we propose a semantic replacement module. During training, a small subset of samples is selected, and a region is randomly chosen based on semantic area size as probability. The selected region is replaced with a region from another sample with the most divergent attributes in the parameter space. This artificial manipulation encourages the model to detect regional discrepancies and thus develop fine-grained retouching capabilities.
Second, although the semantic replacement module facilitates local retouching, the model tends to ignore global coherence, resulting in visually inconsistent and aesthetically unpleasing outputs. This suggests that the model is overly sensitive to the segmentation information contained in the parameter map and not fully utilizing the strong generative capacity of the diffusion prior. To alleviate this, we introduce perturbations to the parameter map along multiple dimensions, such as channel shifts and blurring, thereby enforcing the treatment of the segmentation as soft guidance rather than a rigid structure. This encourages the model to interpret and respond to semantic boundaries implicitly during the generation process. The effectiveness of both modules is further examined in Section 4.3.
@@
To enable fine-grained control over regional image attributes, we integrate ControlNet into the Stable Diffusion framework. To accommodate multi-attribute conditioning, we expand the region-level attribute scores into a multi-channel guidance map $C=\{C^{1},C^{2},\dots,C^{K}\}$ of the same spatial resolution as the segmentation map, where each channel $C^{k}$ encodes the spatial distribution of a specific image attribute (e.g., colorfulness, contrast, color temperature, brightness). Each pixel value in $C^{k}$ reflects the score of its corresponding region in that attribute. By adjusting the values of specific semantic regions within $C$, the model outputs corresponding retouching styles for the associated attributes, while simultaneously maintaining global image aesthetics. The coefficient values are adjusted within the range [1,1][-1,1], where each value corresponds to a visual style learned within the distribution of high-quality images. The specific parameter map injection method and training details are given in the Supplementary Material.
@@ 3.3 Agent
Weak instructions are designed for non-expert users who prefer minimal interaction. In this mode, the agent automatically constructs the multi-channel parameter map $C={C^{1},C^{2},\dots,C^{K}}$ using the midpoint of each image attribute as the default value. While some methods (song2021starenhancer) allow users to provide reference images to indicate their preferred style, we adopt the midpoint-based initialization to reduce user learning costs and improve usability. The parameter map is then further tailored based on the user’s editing history and preferences. This guidance map is then injected into the ControlNet to enable rapid and accurate region-aware retouching, producing visually appealing results that align with the user’s historical aesthetic tendencies, without requiring explicit manual input.
@@ 4.2 Comparisons
To further validate the effectiveness of our approach, we conduct a user study to assess human preferences over PerTouch and other SOTA baselines including DiffRetouch, StarEnhancer, and TSFlow. We randomly select 30 images from the MIT-Adobe FiveK test set and recruit 50 volunteers to participate in the evaluation. Given the original input and retouched results from all methods, participants are asked to choose the result that best aligns with their personal preferences. We calculated the preference percentage for each method per user and summarized the results in the Kernel Density Estimation plot shown in Figure 6. The majority of participants expressed a preference for our method in nearly half of the test cases, significantly outperforming other approaches. This demonstrates the strong ability of our method to generate visually pleasing and user-preferred results.
@@ 5 Conclusion
6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62306153, 62225604), the Natural Science Foundation of Tianjin, China (24JCJQJC00020), the Young Elite Scientists Sponsorship Program by CAST (YESS20240686), the Fundamental Research Funds for the Central Universities (Nankai University, 070-63243143), and Shenzhen Science and Technology Program (JCYJ20240813114237048). This work was also funded by Samsung R&D Institute China-Beijing (SRC-B).
The computational devices are supported by the Supercomputing Center of Nankai University (NKSC).

### v3

@@ Abstract
Image retouching aims to enhance visual quality while aligning with users’ personalized aesthetic preferences. To address the challenge of balancing controllability and subjectivity, we propose a unified diffusion-based image retouching framework called PerTouch. Our method supports semantic-level image retouching while maintaining global aesthetics. Using parameter maps containing attribute values in specific semantic regions as input, PerTouch constructs an explicit parameter-to-image mapping for fine-grained image retouching. To improve semantic boundary perception, we introduce semantic replacement and parameter perturbation mechanisms during training. To connect natural language instructions with visual control, we develop a VLM-driven agent to handle both strong and weak user instructions. Equipped with mechanisms of feedback-driven rethinking and scene-aware memory, PerTouch better aligns with user intent and captures long-term preferences. Extensive experiments demonstrate each component’s effectiveness and the superior performance of PerTouch in personalized image retouching.
Codehttps://github.com/Auroral703/PerTouch
@@ 3.2 Architecture
We observe that directly injecting the parameter maps into the network leads to over-reliance on the information it encodes, which is contrary to our objective. Rather than enforcing strict adherence to the parameter maps, we aim to use them as a soft hint, allowing the diffusion prior to play a central role in generating aesthetically pleasing results. To mitigate the model’s over-reliance on the parameter map, we introduce two mechanisms.
First, we find that the model struggles to perceive region boundaries based on the injected parameter map, often defaulting to global retouching. This is because the input map contains only coarse semantic scores, which are not spatially continuous across regions, while real images often exhibit spatially coherent color transitions. This discrepancy makes it difficult for the model to correlate the parameter maps with the image structure. To address this, we propose a semantic replacement module. During training, a small subset of samples is selected, and a region is randomly chosen based on semantic area size as probability. The selected region is replaced with a region from another sample with the most divergent attributes in the parameter space. This artificial manipulation encourages the model to detect regional discrepancies and thus develop fine-grained retouching capabilities.
Second, although the semantic replacement module facilitates local retouching, the model tends to ignore global coherence, resulting in visually inconsistent and aesthetically unpleasing outputs. This suggests that the model is overly sensitive to the segmentation information contained in the parameter maps and not fully utilizing the strong generative capacity of the diffusion prior. To alleviate this, we introduce perturbations to the parameter maps along multiple dimensions, such as channel shifts and blurring, thereby enforcing the treatment of the segmentation as soft guidance rather than a rigid structure. This encourages the model to interpret and respond to semantic boundaries implicitly during the generation process. The effectiveness of both modules is further examined in Section 4.3.
@@
To enable fine-grained control over regional image attributes, we integrate ControlNet into the Stable Diffusion framework. To accommodate multi-attribute conditioning, we expand the region-level attribute scores into a multi-channel guidance map $C=\{C^{1},C^{2},\dots,C^{K}\}$ of the same spatial resolution as the segmentation map, where each channel $C^{k}$ encodes the spatial distribution of a specific image attribute (e.g., colorfulness, contrast, color temperature, brightness). Each pixel value in $C^{k}$ reflects the score of its corresponding region in that attribute. By adjusting the values of specific semantic regions within $C$, the model outputs corresponding retouching styles for the associated attributes, while simultaneously maintaining global image aesthetics. The coefficient values are adjusted within the range [1,1][-1,1], where each value corresponds to a visual style learned within the distribution of high-quality images. The specific parameter maps injection method and training details are given in the Supplementary Material.
@@ 3.3 Agent
Weak instructions are designed for non-expert users who prefer minimal interaction. In this mode, the agent automatically constructs the multi-channel parameter maps $C={C^{1},C^{2},\dots,C^{K}}$ using the midpoint of each image attribute as the default value. While some methods (song2021starenhancer) allow users to provide reference images to indicate their preferred style, we adopt the midpoint-based initialization to reduce user learning costs and improve usability. The parameter maps are then further tailored based on the user’s editing history and preferences. This guidance map is then injected into the ControlNet to enable rapid and accurate region-aware retouching, producing visually appealing results that align with the user’s historical aesthetic tendencies, without requiring explicit manual input.
@@ 4.2 Comparisons
To further validate the effectiveness of our approach, we conduct a user study to assess human preferences over PerTouch and other SOTA baselines including DiffRetouch, StarEnhancer, TSFlow, and JarvisArt. We randomly select 30 images from the MIT-Adobe FiveK test set and recruit 50 volunteers to participate in the evaluation. Given the original input and retouched results from all methods, participants are asked to choose the result that best aligns with their personal preferences. We calculated the preference percentage for each method per user and summarized the results in the Kernel Density Estimation plot shown in Figure 6. The majority of participants expressed a preference for our method in nearly half of the test cases, significantly outperforming other approaches. This demonstrates the strong ability of our method to generate visually pleasing and user-preferred results.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62306153, 62225604), the Natural Science Foundation of Tianjin, China (24JCJQJC00020), the Young Elite Scientists Sponsorship Program by CAST (YESS20240686), the Fundamental Research Funds for the Central Universities (Nankai University, 070-63243143), and Shenzhen Science and Technology Program (JCYJ20240813114237048). This work was also funded by Samsung R&D Institute China-Beijing (SRC-B). The computational devices is supported by the Supercomputing Center of Nankai University (NKSC).